The prices of CD players are normally distributed, with a standard deviation of $12. A random sample of 16
CD player prices is drawn. A confidence interval for the mean price shows a $9.30 margin of error. What was
the level of confidence?
Answer #1

Given that, standard deviation (sigma) = $12

sample size ( n ) = 16

Margin of error ( E ) = $9.30

where,

E = Z_{alpha/2} * rac { sigma}{sqrt { n }}

12 => 9.30 = Za/2 * To

12 >9.30 = Za/2 * 4

9.30 3 * Zo/2

9.30

Za/23.10

Using standard normal z-table we get,

P(Z < 3.10) = 0.9990

Therefore, we get, a/2-1-0.9990 0.001

a-2 * 0.001 0.002

Hence, level of confidence ( c ) = 1-a1- 0.0020.998

Level of confidence =  99.8%